Minicase 3 The Mallory Corporation On December 31, 2006, the Mallory Corporation had the following activity in its fixed assets record. MALLORY CORPORATION - FIXED ASSETS Equipment Cost Salvage Life Method of Depreciation Machine 1 $65,000 $5,000 5 DDB purchased 1/1/2006 Building #3 $900,000 not including land $50,000 25 S/L purchased 6/30/2006 Mine 316 $1,000,000 $0 1,000,000 tons 30,000 tons extracted. Mine purchased 1/1/2006 Patent $50,000 0 17 Purchased 1/1/2006 Truck 1 $35,000 $3,000 200,000 miles Units of production: total miles depreciated to date are 60,000 as of January 1, 2006. Miles this year 30,000 REQUIRED: • Compute the depletion, amortization, and depreciation expense on December 31, 2006 for each asset listed above • Record the depreciation journal entries for the assets above • Suppose that Machine 1 was sold for $40,000 on 12/31/2008, record the entry • Suppose that the corporation spent $20,000 in 2006 to defend the patent. Record the entry. • Financial Reporting on Fixed Assets:  Prepare a partial balance sheet statement for Mallory Corporation showing Fixed and Intangible assets
Answer #1

ANSWER:-

1).

Computation of depletion, amortization, and Depreciation

Machine 1 (Depreciation) DDB = 2 * Straight line * Book value of beginning of the year

= 2 * 20% (1/5) * 65,000

= 26,000

Building

= (90,000 -50,000) * 25

= 1,600

Mine 316 tons (depletion) = (Cost - Salvage value) / Estimated no of units * No of units extracted

= (1,000,000-0) / 1,000,000*30,000

   = 30,000

Mine 682 tons = (500,000/40,000)*6,000

= 75,000

Patents amortization = 50,000 / 17

= 2941.1765

Truck Depreciation = (35,000 - 3,000) / 200,000 * 60,000

= 9,600

Truck 2 = (50,000 - 5,000) / 150,000 * 15,000

= 4,500

Truck 3 = (75,000 - 10,000) / 180,000 * 200,000

= 72,222.2222

   Machine 2    = (100,000 - 5,000) / 10

= 9,500

2).

Cash 50,000

Loss on sale 45,000

Machine 2 90,500

3).

Building = (90,000 - 5,000) / 30

= 1333.3333

Depreciation 1333.3333

Building 1333.3333

4).

Patent 20,000

Cash 20,000

-----------------------------------------------------------------

If you have any queries please ask me in the comment i am here for help you. Please do not direct thumbs down just ask if you have any query. And if you like my work then please appreciates with up vote.

                                                     ************THANK YOU************